Consider a hypothesis economy described by the following equation 

C=100

I=1200

X-1110

M=200+0.25Y

T=250+0.3Y

Required 

1.Compute the equilibrium level of the national income

2.The level of consumption income after tax and net exports that corresponds to the equilibrium level

 national income


1
Expert's answer
2021-06-24T12:43:24-0400

1:

National income"(Y)=C+I+G+X-M\\Y=100+1200+0+1110-(200+0.25Y)\\Y=1300+1110-200-0.25Y\\Y+0.25Y=2410-200\\1.25Y=2210\\Y=\frac{2210}{1.25}\\Y=1768"


2:

Level of consumption income after tax(Yd)= National income (Y)-Tax(T)

"Y=1768\\T=250+0.3Y=250+0.3(1768)\\T=780.4\\Y_D=Y-T=1768-780.4\\Y_D=987.6"

Net exports=X-M

"X=1110\\M=200+0.25Y\\M=200+0.25×1768\\M=200+442\\M=642"

Net exports"=1110-642=468"

The level of consumption income after tax is 987.6 and net exports is 468